# A Reconfigurable Time-Domain In-Memory Computing Macro using FeFET-Based CAM with Multilevel Delay Calibration in 28 nm CMOS

### v2

@@ I Introduction
TD-nvIMC offers improved signal margin over voltage-domain approaches, particularly when performing a large number of accumulations, as illustrated in Fig. 1(b). In TD-nvIMC, decreasing the time resolution or least significant bit duration ($T_{LSB}$) improves both latency and throughput. However, this reduction also narrows the timing margin, making the system more susceptible to noise and variations. To sustain high performance while preserving signal integrity, calibration techniques are essential to compensate for device mismatch and ensure reliable operation at fine time resolutions [see Fig. 1(b)]. As qualitatively illustrated in Fig. 1(c), TD-nvIMC achieves a favorable trade-off across key performance metrics, such as area efficiency, scalability, digital compatibility and energy efficiency, compared to digital and voltage-domain IMC architectures.
@@ III-A Memory Array Design, Write-Disturb Prevention
To erase a cell to HVT, -4 V is applied to the WL of the selected cell, while the BL, SL, and BuL are held at 0 V. To prevent write disturbance, the other cells in the same column are isolated by biasing the BuL to -2 V. This negative bulk bias applied to unselected rows suppresses unintended polarization switching during write pulses for the cells in the same column. Triple-well isolation ensures that each row has a dedicated bulk node, enabling independent control of the BuL signals per row. Jiang et al[28] showed that this scheme is effective at the device level, demonstrating that bulk biasing can prevent switching in individual FeFETs. In this work, we validate this approach at the array level through experimental measurements (see Fig 6), confirming that only the selected cell switches to HVT while all unselected cells remain undisturbed. To the best of our knowledge, this is the first array-level validation of bulk-assisted write-disturb prevention in FeFET-based C-AND arrays.
Figure 5: C-AND schematic and write scheme. First, all cells in the selected column are programmed to LVT. Erase is performed by selecting the target cell (WL=4VWL=-4~V) and setting BuL=0VBuL=0~V. Write-disturb is prevented on adjacent cells in the column with BuL=2VBuL=-2~V.
Figure 6: C-AND schematic and write scheme. First, all cells in the selected column are programmed to LVT. Erase is performed by selecting the target cell (WL=4VWL=-4~V) and setting BuL=0VBuL=0~V. Write-disturb is prevented on adjacent cells in the column with BuL=2VBuL=-2~V.
@@ III-D DE and CSI
$R_{CAM}=\frac{L_{FeFET}/W_{FeFET}}{k_{FeFET}(V_{WL}-V_{T,FeFET})}\parallel\frac{L_{FeFET}/W_{FeFET}}{k_{FeFET}(V_{\overline{WL}}-V_{T,\overline{FeFET}})}.$ (1)
@@ IV-A XOR-MAC
The XOR-MAC operation is performed by grounding the selected BL, while setting all other BLs to Hi-Z (Fig. 8). The activation inputs $X_{i}$ are applied to the WL and their complements ($\overline{X_{i}}$) to $\overline{\text{WL}}$. Logic ‘1’ is set to $V_{H}$, which is set during the calibration phase (see Sec. V). A match, $X_{i}=W_{i,j}$, leads to a fast discharge path through a LVT FeFET with a high gate voltage, resulting in $t_{dL}$ delay, while $X_{i}\neq W_{i,j}$ disables the fast discharge path either because the FeFET being in HVT state, or due to the gate voltage being low (0 V). In this case, the DE discharges only through the leaker, producing $t_{dH}$ delay. This conditional delay behavior stems from the effective resistance at the tail of the CSI, modeled as $R_{CAM}\parallel R_{leaker}$, as described by expressions  (1)–(3).
@@ IV-C Boolean Logic
In addition to MAC operations, the reconfigurable macro supports in-memory logic operations such as OR and AND, as well as full adder operations, performed directly on the stored FeFET states. These operations are executed per row, using a subset of $k$ selected columns out of total $M$. A row is selected by grounding the corresponding BL, while all other BLs are set to Hi-Z, as in the MAC mode. The participating columns are selected by applying $V_{H}$ to their WLs, while grounding all other WLs and all $\overline{WL}$ signals (including for selected cells). This ensures that the DEs of the unselected cells remain in their high-delay state, unaffected by the stored memory (FeFET state).
@@ V Calibration Scheme
VI Experimental Results of MAC and Logic Operations
@@
Figure 13: Experimental setup. Two MCPs with 15 probes each are used for on-die probing. The IO driver output is sampled using a high-speed oscilloscope with $50~\Omega$ input. The control signals and input pulse are generated by AWGs and power supplies are used for $V_{DD}$ and bias signals such as $V_{leak}$, step, and shift.
The fabricated proof-of-concept (PoC) includes a $\times3$ CAM array, a 3-stage DE chain, a 2-bit TDC, a reference delay line, I/O driver, decoder, and testing circuits, and is powered by a 0.85-V supply ($V_{DD}$). An optical micrograph of the fabricated PoC is shown in Fig. 12. The PoC macro area is $17.6\times 30.7~\mu m^{2}$. Note that the I/O occupies a significant area while the CAM array, DE chain, and TDC occupy a relatively compact footprint. The experimental setup is illustrated in Fig. 13. On-die measurements are performed using two multicontact probes (MCPs) with 15 probes each. The I/O driver output is sampled using a high-speed oscilloscope with $50~\Omega$ input. The power supply, control signals, and input pulse are generated by an arbitrary waveform generator (AWG) and DC supplies.
@@
XOR-MAC is experimentally validated, covering all $2^{6}=64$ input combinations of three binary activations-weights pair. Each configurations is loaded into the CAM array and triggered through the DE chain. The resulting output pulse (delay) is sampled at $DEC_{O}$ and digitized using the 2-bit TDC, which maps the delay to a corresponding $TDC_{O}$ code. The observed step size between successive MAC results is 1.3 ns, with four distinct levels clearly distinguishable as shown in Fig. 14. This confirms the ability of the macro to resolve the full dynamic range of XOR-based MAC operations.

### v3

@@ I Introduction
TD-nvIMC offers improved signal margin over voltage-domain approaches, particularly when performing a large number of accumulations, as illustrated in Fig. 1(b). In TD-nvIMC, decreasing the time resolution or least significant bit duration ($T_{LSB}$) improves both latency and throughput. However, this reduction also narrows the timing margin, making the system more susceptible to noise and variations, approaching the reduced signal margins typical of voltage-domain implementations. To sustain high performance while preserving signal integrity, calibration techniques are essential to compensate for device mismatch and ensure reliable operation at fine time resolutions [see Fig. 1(b)]. As qualitatively illustrated in Fig. 1(c), TD-nvIMC achieves a favorable trade-off across key performance metrics, such as area efficiency, scalability, digital compatibility and energy efficiency, compared to digital and voltage-domain IMC architectures.
@@ III-A Memory Array Design, Write-Disturb Prevention
To erase a cell to HVT, -4 V is applied to the WL of the selected cell, while the BL, SL, and BuL are held at 0 V. To prevent write disturbance, the other cells in the same column are isolated by biasing the BuL to -2 V. This negative bulk bias applied to unselected rows suppresses unintended polarization switching during write pulses for the cells in the same column. Triple-well isolation ensures that each row has a dedicated bulk node, enabling independent control of the BuL signals per row. Jiang et al[11] showed that this scheme is effective at the device level, demonstrating that bulk biasing can prevent switching in individual FeFETs. In this work, we validate this approach at the array level through experimental measurements (see Fig. 6), confirming that only the selected cell switches to HVT while all unselected cells remain undisturbed. To the best of our knowledge, this is the first array-level validation of bulk-assisted write-disturb prevention in FeFET-based C-AND arrays.
Figure 5: C-AND schematic and write scheme. (a) First, all cells in the selected column are programmed to LVT. (b) Erase is performed by selecting the target cell (WL=4VWL=-4~V) and setting BuL=0VBuL=0~V. Write-disturb is prevented on adjacent cells in the column with BuL=2VBuL=-2~V.
Figure 6: Array-level write-disturb prevention using bulk biasing. Measured $I_{BL}$$V_{WL}$ characteristics for a selected cell, half-selected cells sharing the same BuL or WL, and an unselected cell. Only the selected cell switches to HVT after erase, while all other cases show no disturb.
@@ III-D DE and CSI
$R_{CAM}=\frac{L_{FeFET}/W_{FeFET}}{k_{FeFET}(V_{WL}-V_{T,FeFET})}\parallel\frac{L_{FeFET}/W_{FeFET}}{k_{FeFET}(V_{\overline{WL}}-V_{T,\overline{FeFET}})}{\color[rgb]{0,0,0},}$ (1)
@@ IV-A XOR-MAC
The XOR-MAC operation is performed by grounding the selected BL, while setting all other BLs to Hi-Z (Fig. 8). The activation inputs $X_{i}$ are applied to the WL and their complements ($\overline{X_{i}}$) to $\overline{\text{WL}}$. Logic ‘1’ is set to $V_{H}$, which is set during the calibration phase (see Sec. V). A match, $X_{i}=W_{i,j}$, leads to a fast discharge path through a LVT FeFET with a high gate voltage, resulting in $t_{dL}$ delay, while $X_{i}\neq W_{i,j}$ disables the fast discharge path either becaue the FeFET being in HVT state, or due to the gate voltage being low (0 V). In this case, the DE discharges only through the leaker, producing $t_{dH}$ delay. This conditional delay behavior stems from the effective resistance at the tail of the CSI, modeled as $R_{CAM}\parallel R_{leaker}$, as described by expressions  (1)–(3).
@@ IV-C Boolean Logic
In addition to MAC operations, the reconfigurable macro supports in-memory logic operations such as OR and AND, as well as full adder operations, performed directly on the stored FeFET states. These operations are executed per row, using a subset of $k$ selected columns out of total $M$. A row is selected by grounding the corresponding BL, while all other BLs are set to Hi-Z, as in the MAC mode. The participating columns are selected by applying $V_{H}$ to their WLs, while grounding all other WLs and all $\overline{\text{WL}}$ signals (including for selected cells). This ensures that the DEs of the unselected cells remain in their high-delay state, unaffected by the stored memory (FeFET state).
@@ V Calibration Scheme
The impact of MLS calibration on computation correctness is reflected in the timing margin between adjacent time-domain outcomes. Reliable digitization requires that the accumulated delay uncertainty remains well below the measured separation $\Delta s$ between neighboring levels. The total timing uncertainty can be written as
$\sigma_{T}^{2}=k\,\sigma_{dL}^{2}+(N-k)\,\sigma_{dH}^{2}+\sigma_{\mathrm{jit}}^{2}+\sigma_{\mathrm{TDC}}^{2},$ (5)
where $N$ is the number of DEs in the chain and $k$ the number of DEs contributing the short-delay state. The terms $\sigma_{dL}$ and $\sigma_{dH}$ denote the stage-to-stage dispersion of the short- and long-delay states, respectively, while $\sigma_{\mathrm{jit}}$ and $\sigma_{\mathrm{TDC}}$ represent other jitter sources (e.g., supply noise, thermal noise) and TDC uncertainty. Without calibration, $\sigma_{dL}$ is dominated by intrinsic D2D mismatch, which accumulates with $k$ and reduces the margin between adjacent delay levels. With MLS calibration, the residual D2D dispersion of $t_{dL}$ is bounded by the tuning resolution,
$\sigma_{dL,\mathrm{D2D}}^{\mathrm{calibrated}}\leq\frac{\Delta{t}}{\sqrt{12}},\qquad\Delta{t}\leq 100~\mathrm{ps},$ (6)
thereby replacing the process-limited D2D spread with a resolution-limited bound. In contrast, $t_{dH}$ is stabilized by the leaker path, which clamps its delay and limits sensitivity to FeFET off-state variability.
Assuming midpoint decision thresholds and Gaussian timing uncertainty, the probability of misclassification between adjacent outcomes is
$P_{\mathrm{err}}=Q\!\left(\frac{\Delta s}{2\sigma_{T}}\right),$ (7)
where $Q(\cdot)$ is the complementary Gaussian cumulative distribution function. For a fixed $\Delta s$, the impact of calibration on computation correctness is captured by the ratio of timing uncertainties,
$\frac{\sigma_{T}^{\mathrm{uncalibrated}}}{\sigma_{T}^{\mathrm{calibrated}}}\approx\sqrt{\frac{k\,(\sigma_{dL,\mathrm{D2D}}^{\mathrm{uncalibrated}})^{2}+\Phi}{k\,(\Delta{t}/\sqrt{12})^{2}+\Phi}},$ (8)
where $\Phi=(N-k)\sigma_{dH}^{2}+\sigma_{\mathrm{jit}}^{2}+\sigma_{\mathrm{TDC}}^{2}$ groups all non-calibrated contributions. Since $\sigma_{dL,\mathrm{D2D}}^{\mathrm{uncalibrated}}$ dominates prior to calibration, MLS tuning increases the timing margin by directly suppressing the leading D2D term. This margin restoration exponentially reduces $P_{\mathrm{err}}$ via (7), explaining the experimentally observed separation of MAC and logic delay levels and establishing the accuracy gain required for reliable operation under variability when scaling to larger arrays.
@@
VI Experimental Results
VI-A Experimental Results of MAC and Logic Operations
@@
Figure 13: Experimental setup. Two MCPs with 15 probes each are used for on-die probing. The IO driver output is sampled using a high-speed oscilloscope with $50~\Omega$ input. The control signals and input pulse are generated by AWGs, write pulses by the WGFMU, and power supplies are used for $V_{DD}$ and bias signals such as $V_{leak}$, step, and shift.
The fabricated proof-of-concept (PoC) includes a $\times3$ CAM array, a 3-stage DE chain, a 2-bit TDC, a reference delay line, I/O driver, decoder, and testing circuits, and is powered by a 0.85-V supply ($V_{DD}$). An optical micrograph of the fabricated PoC is shown in Fig. 12. The PoC macro area is $17.6\times 30.7~\mu m^{2}$. Note that the I/O occupies a significant area while the CAM array, DE chain, and TDC occupy a relatively compact footprint. The experimental setup is illustrated in Fig. 13. On-die measurements are performed using two multicontact probes (MCPs) with 15 probes each. The I/O driver output is sampled using a high-speed oscilloscope with $50~\Omega$ input. DC supplies provide the supply voltage ($V_{DD}$) and bias signals, while the control signals and input pulse are generated by an arbitrary waveform generator (AWG). Furthermore, a waveform generator/fast measurement unit (WGFMU) is used to generate the program and erase pulses for the FeFETs and to perform current measurements for device characterization.
@@
XOR-MAC is experimentally validated, covering all $2^{6}=64$ input combinations of three binary activations-weights pair. Each configuration is loaded into the CAM array and triggered through the DE chain. The resulting output pulse (delay) is sampled at $DEC_{O}$ and digitized using the 2-bit TDC, which maps the delay to a corresponding $TDC_{O}$ code. The observed step size between successive MAC results is 1.3 ns, with four distinct levels clearly distinguishable as shown in Fig. 14. This confirms the ability of the macro to resolve the full dynamic range of XOR-based MAC operations.
@@
VI-B FeFET Retention and Endurance Characterization
The proposed TD-nvIMC macro relies on stable FeFET $V_{T}$ states to support the MLS-based calibration in Section III-B. Prior work on HfO2-based FeFET arrays in 28-nm CMOS reports stable retention up to $10^{4}$ s at 85C and endurance on the order of $10^{3}$$10^{4}$ full-swing cycling before stronger degradation at higher stress [30]. In Addition, multilevel FeFET studies report that intermediate states can be as stable as the saturated states [17, 3]. Here, retention was evaluated by tuning devices to multiple MLS and monitoring $I_{BL}$ over $10^{5}$ s ($\approx 28$ h) at room temperature using read pulses with $V_{WL}=0.6$ V and $V_{BL}=0.8$ V. As shown in Fig. 17(a), the measured currents exhibit negligible drift across the tested states. This stability is most critical for the LVT-programmed levels that directly determine $t_{dL}$, while potential drift in the HVT state is further desensitized by the parallel leaker path (Fig. 4), which clamps $t_{dH}$ against minor off-state variations.
Endurance is characterized under repeated full program/erase (P/E) cycling to provide a conservative reliability bound under worst-case switching stress. The measured $I_{BL}$ as a function of P/E cycles using full-swing pulses is shown in Fig. 17(b). The read current remains consistent over the initial cycling regime (up to $\sim 10^{3}$ cycles), while a gradual shift is observed as cycling progresses into the several-thousand-cycle range, which is expected under aggressive full-swing P/E stress. Importantly, the proposed calibration flow is less stressful than this measurement: it starts from a single program operation and then applies controlled bulk-assisted partial erase adjustments. Prior work has reported that partial switching can improve endurance compared to repeated full-swing cycling [9], consistent with reports that bipolar switching accelerates degradation through charge trapping and interface degradation under repeated program/erase switching stress [12]. Therefore, the effective endurance margin in normal calibrated operation is expected to be higher than what is implied by worst-case full-swing cycling.
Figure 17: FeFET reliability characterization. (a) Retention: measured $I_{BL}$ over time for multiple MLS, showing stable readout up to $10^{5}$ s. (b) Endurance: measured $I_{BL}$ versus full-swing program/erase cycles.